OpenStudy (anonymous):

The height of an equilateral triangular prism increases by three units. The new volume is more than the original by how much? three more than the the area of the base three more than the length of the height three times the area of the base three times the length of the height

OpenStudy (experimentx):

the volume is given by = area of base x height now if height is increased by 3 times, then new volume = area of base x 3 height = 3 (area of base x height) = 3 old volume

OpenStudy (anonymous):

thank you for the explanation, but can you tell me if the answer is either A.B.C. or D?

OpenStudy (anonymous):

??

OpenStudy (experimentx):

which do you think??

OpenStudy (anonymous):

@KailynBarratt using the information he has given you, can you?

OpenStudy (anonymous):

hmmm, C?

OpenStudy (anonymous):

well, those are the answers i have to choose from on the practice sheet.. so it has to be one of those choices. :/

OpenStudy (anonymous):

nope, copied and pasted it lol. it's ok, i'll call me teacher and ask. i appreciate the help though!

OpenStudy (experimentx):

I might not be understanding the what exactly the question is implying, the new volume will be three times bigger than the original one. you can achieve this if you increase the base by 3 times, the last option would also have same effect. I seriously doubt that the options are correct though. I suggest you consult the question maker.

Directrix (directrix):

three times the area of the base ---------------------------- Original Prism Volume = Bh where B is the area of the base and h is the height of the prism. New Prism Volume = B(h+3) where the height of the original prism has been increased by 3.. V = B(h+3) V = Bh + 3B ---> compare to V = Bh